OpenStudy (lisa123):

f(x)= x^3-1/ x^2+x+1. Find the domain, zeros, holes, and asymptotes

OpenStudy (anonymous):

do you know how to factor the difference of two cubes?

OpenStudy (lisa123):

(x-1)(x^2+x+1)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now cancel and you are done

OpenStudy (lisa123):

I don't know how to factor the denominator

OpenStudy (anonymous):

take a look at your numerator in factored form

OpenStudy (lisa123):

ohhhhh wowww

OpenStudy (anonymous):

\[\frac{(x-1)(x^2+x+1)}{x^2+x+1}\] no need to factor in fact you cannot factor further, \(x^2+x+1\) has no zeros

OpenStudy (lisa123):

so what is the domain?

OpenStudy (anonymous):

since the denominator has no zeros, the domain is all real numbers

OpenStudy (anonymous):

normally you would say "all real numbers except for the zeros in the denominator" but this denominator doesn't have any is it clear why?

OpenStudy (lisa123):

yeah b/c it can only be factored with imaginary numbers

OpenStudy (anonymous):

right

OpenStudy (lisa123):

what are the asymptotes....I know its oblique

OpenStudy (anonymous):

hold the phone

OpenStudy (anonymous):

the whole thing is just the line \(f(x)=x-1\)

OpenStudy (anonymous):

it is a line it does not have asymptotes

OpenStudy (lisa123):

I thought you looked for the horizontal asymptote based on the exponents

OpenStudy (anonymous):

normally you would

OpenStudy (anonymous):

but in this case the entire denominator cancels

OpenStudy (anonymous):

\[\frac{(x-1)(x^2+x+1)}{x^2+x+1}=x-1\]

OpenStudy (lisa123):

oh alright thank you!